OpenStudy (anonymous):

@SithsAndGiggles @zepdrix Let k be a constant. The solution f the equation 2x+7=kx-4 is A) 4/(k-2) B) (2x-11)/k C) (kx-11)/2 D) 11/(k-2) E) none of these

OpenStudy (anonymous):

2x+7=kx-4 You have to solve this for x. 7+4=kx-2x 11=x(k-2) x=11/(k-2)